Introduction to and Resources for TEX and L^{A}TEX

L^{A}TEX (more generally, the TEX system by Donald Knuth, on which L^{A}TEX is built) is a powerful software system for typesetting technical documents. It’s a markup language, like html, meaning you write a .tex file in plain text, compile it (for example, with the command latex file.tex in Linux), and the result is a .dvi (DeVice Independent) file, which can be printed or converted to a .ps or .pdf format. It is used almost universally by mathematicians and other scientists. The language is relatively easy to learn, in my opinion, and once you’ve got the hang of it, typing math can be as fast or faster than writing (I have used L^{A}TEX to take notes in some of my classes, admittedly ones with few drawings). A basic example: to get the expression f(x) = \frac{1}{2}x^{2}, you write $f(x) = \frac\{1\}\{2\} x^2$ (the dollar signs delimit a “math environment”), and L^{A}TEX takes care of the rest. Creating things like matrices takes a little more work, but not much. Below are some resources, if you’d like to learn about using TEX; the first website includes some information and links about how to get L^{A}TEX on your computer. It’s usually included with Linux distributions. Oh, and all the software is free, and most is licenced under the GNU project.

   • Various Documentation: http://www.ling.upenn.edu/advice/latex.html

• “L^{A}TEX: from quick and dirty to style and finesse”:

     http://www.sci.usq.edu.au/staff/robertsa/LaTeX/latexintro.html

• MiKTEX: Windows implementation of TEX: http://www.miktex.org/

   • TEX frequently asked questions: http://www.tex.ac.uk/cgi-bin/texfaq2html
   • L^{A}TEX help: http://www.emerson.emory.edu/services/latex/latex toc.html

• CTAN: the Comprehensive TEX Archive Network: http://www.ctan.org/

The book I use the most: Math Into L^{A}TEX, Third Edition, by George Grätzer. The first chapter (including

“Typing your first article”) is available at http://www.ctan.org/tex-archive/info/mil/mil.pdf.

Some other related software I’ve found useful:

• , a graphics system that will draw lines, points, curves, etc. Finding clear documentation is difficult: the Page has links to documents written by the creator, John Hobby, but I usually use the examples created by Vincent Zoonekynd to figure out what I need to do.

     http://cm.bell-labs.com/who/hobby/MetaPost.html
     http://zoonek.free.fr/LaTeX/Metapost/metapost.html

• ArabTEX, for typesetting Arabic, Hebrew, and other Middle Eastern languages, by Klaus Legally.

     http://www.informatik.uni-stuttgart.de/ifi/bs/research/arab_e.html

• Lilypond, a music notation system by Han-Wen Nienhuys and Jan Nieuwenhuizen. It uses TEX internally (I think) for some typesetting commands, but it has its own syntax. The results really are beautiful, if you write music and like to have it look good. I’ve also read that it’s possible to incorporate Lilypond source into a L^{A}TEX file and thus have music examples in your .dvi files, but I haven’t managed to do it, yet.

     http://www.lilypond.org/web/